# Riemannian Metric Learning: Closer to You than You Imagine

### v1

@@ Abstract
To recount the story from the beginning, let us begin with the principles of metric learning. Metric learning [20] is an integral, though often implicit, aspect of machine learning. Whether consciously or not, most practitioners in the field are already engaging with it—or will eventually. The concept has a long history, dating back to early methods like the k-nearest neighbors (k-NN) classifier [42, 164, 64]. As the term ‘nearest’ suggests, a measure of distance is at the heart of k-NN algorithms as depicted in Figure 1. Without prior knowledge, the distance is often selected as the Euclidean distance. However, this choice is uninformative of the input and output distribution. That is why Atkeson et al. [10] suggested to learn the distance by learning a symmetric positive definite matrix $G\in\mathbb{R}^{d\times d}$ such that $\mathbf{d}_{G}(x,y)=\sqrt{(y-x)^{\top}G(y-x)}$. It turns out that learning $G=L^{\top}L$ is equivalent to learning a linear transformation $x\mapsto Lx$ of the input space as $\mathbf{d}_{G}=\mathbf{d}(L\cdot,L\cdot)$. The matrix $G$ can be learned, for example, in an unsupervised manner using Principal Component Analysis (PCA) or in a supervised manner using Linear Discriminant Analysis (LDA). Learning $G$ is valuable for two key reasons: First, it assigns global relative weights to the different coordinates when computing distances. Second, it allows us to account for global correlations in the dataset. However, $G$ needs to vary to account for heterogeneous local geometries within the dataset as shown in Figure 1. This linear framework was then extended with non-linear transformations $x\mapsto\phi(x)$ by using kernels or neural networks [198, 52, 66] such that the distance is $\mathbf{d}(\phi(\cdot),\phi(\cdot))$, which reduces to learning a feature function $\phi$ as already pointed out.
@@
Going one step further, metric learning can be embedded in a more general framework called Riemannian metric learning [143], in the same way that Galilean relativity is embedded by General relativity [57]. In Riemannian metric learning, the input space $(\mathcal{X},\mathfrak{g})$ is a Riemannian manifold, i.e. a kind of smooth geometric shape (see Figure 2) with a Riemannian metric $\mathfrak{g}\colon\mathcal{X}\to\mathbb{R}^{d\times d}$ that allows to locally measure lengths and angles via a matrix field. From this local geometric structure, one can define a global distance between any $x,y\in\mathcal{X}$ as
@@
Minimizing curves of (1.1) are called metric geodesics and can be understood as generalizations of straight lines to manifolds as illustrated in Figure 1. If the Riemannian metric is constant and the manifold is the Euclidean space, then geodesics are straight lines and we recover the distance metric learning framework from above. However, Riemannian metric learning is not just a generalization of distance metric learning with non-linear transformations; it also gives new mathematical objects useful for modeling. For instance, geodesics can express trajectories for longitudinal modeling and interpolation [123, 109, 100], parallel transport (defined later) can be used in transfer learning [172, 73], and the metric volume can be used for modeling probability distribution and sampling [44, 47]. All these mathematical objects are related to each other and have been extensively studied in Riemannian Geometry [50, 166]. Riemannian metric learning leverages the wealth of this literature to offer both flexible and interpretable modeling tools.
@@
First, we recall preliminary concepts on the essential tools and properties of Riemannian geometry in Section 2; readers familiar with these notions may safely skip this section. At the same time, we introduce the main manifolds relevant to the data used in Riemannian metric learning in Section 2.1, along with the manifold of Riemannian metrics, to provide a broader perspective on the subject. Second, the formal optimization problem associated with Riemannian metric learning is presented in Section 3.1 to define the topic rigorously. This is followed by an overview of its numerous applications in Section 3.2, highlighting the practical benefits and key techniques of Riemannian formalism. Third, we discuss how to perform Riemannian metric learning in Section 4, classifying the methods based on their metric parameterization (Section 4.1) and associated computational approaches (Section 4.1.3), their objectives (Section 4.2), and their optimization strategies (Section 4.3). This classification helps clarify the various technical challenges that arise when designing Riemannian metric learning methods. Finally, before concluding in Section 6, we outline key research directions in Riemannian metric learning in Section 5, which may inspire both theorists and practitioners in the applied mathematics community.
@@ Acknowledgments
[…]
@@
A differentiable manifold $\mathcal{M}\subset\mathbb{R}^{d_{a}}$ is formally defined in Appendix A or in [112, Chapter 2], where $d_{a}$ denotes the dimension of the ambient space. For a first reading, and to simplify notation [^1] All concepts introduced below are compatible with changes of parametrization in overlapping regions and can therefore be generalized to manifolds requiring multiple parametrizations., a differentiable manifold of dimension $d<\infty$ can be understood as the image of a global, surjective, and smooth parametrization $\psi\colon U\subset\mathbb{R}^{d}\to\mathcal{M}$, where $U$ is an open subset of $\mathbb{R}^{d}$. In general, we would need more than one parametrization and all concepts presented hereafter generalize to this scenario. Readers unfamiliar with this concept may safely think of a differentiable manifold as a smoothly curved, nonlinear object, such as a curved surface in three-dimensional space as depicted in Figure 2. Note that in all the following a function $f\colon\mathcal{M}\to\mathcal{Y}\subset\mathbb{R}^{m}$ is called smooth or infinitely differentiable if $f\circ\psi\colon U\to\mathcal{Y}$ is so, which offers a valid definition since $U$ is an open set of $\mathbb{R}^{d}$.
@@
The dimension of the manifold is thus given by the dimension of possible local displacements in each point. In the case of two-dimensional surfaces, the tangent space is simply the plane that is tangent to the surface at a given point as illustrated in Figure 2.
@@
More manifold examples and insights related to computations and modeling will be given in Section 2.1.
@@
This path length is invariant to reparametrization of the curve [112, Lemma 6.1]. Furthermore, the path energy is defined as
@@
and equality holds if and only if $\mathfrak{g}_{\gamma(t)}(\dot{\gamma}(t),\dot{\gamma}(t))$ is constant along the path, which is the case for geodesics as we will see later. Given the length of paths, we can define the distance between two arbitrary points as the shortest possible length of a path connecting them, which leads to the definition of a Riemannian distance [112, Lemma 6.2]:
[…]
@@
Shortest paths can have different shapes depending on the metric as illustrated in Figure 3.
@@
Note that not all distances on manifolds are Riemannian distances. For instance, for any injective function $f\colon\mathcal{M}\to\mathcal{Y}$, $\mathbf{d}_{f}(\cdot,\cdot)\triangleq||f(\cdot)-f(\cdot)||$ is a valid distance on $\mathcal{M}$. However, a metric $\mathfrak{g}$ such that $\mathbf{d}_{f}=\mathbf{d}_{\mathfrak{g}}$ does not exist necessarily. For example, if $f=\mathrm{Id}$ maps the unit circle $\mathsf{S}^{1}$ to $\mathbb{R}^{2}$, the induced distance $\mathbf{d}_{f}$ will approximate the intrinsic distance on the circle only when $x$ and $y$ are sufficiently close. For distant points on the circle, the map-induced distance will underestimate the true geodesic distance.
@@
[…]
@@
We define a covariant derivative $\nabla$ as a map $C^{\infty}(\mathcal{M},T\mathcal{M})\times C^{\infty}(\mathcal{M},T\mathcal{M})\to C^{\infty}(\mathcal{M},T\mathcal{M})$ which satisfies intuitive derivation rules as exposed in the following example or in Appendix A given in Appendix A or in [112, Chapter 4, p.50]. Intuitively, $\nabla_{W}V$ describes how the vector field $V$ changes when moving in direction $W$. In this way, a covariant derivative $\nabla$ of a vector field $V=\sum_{i=1}^{d}V_{i}\partial_{i}\psi$ in direction of $W=\sum_{i=1}^{d}W_{i}\partial_{i}\psi$ denoted by $\nabla_{W}V$ can be decomposed as
@@
for any $x\in\mathcal{M}$ and $Z,V,W\in C^{\infty}(\mathcal{M})$. This defines the Levi-Cevita covariant derivative $\nabla$ also called the Levi-Cevita connection [112, Theorem 5.4]. The following property gives a concrete expression.
[…]
Let $\mathfrak{g}^{-1}=(\mathfrak{g}^{ij})_{(i,j)\in~[d]^{2}}$ be the inverse of $\mathfrak{g}$ as before. Then, the Levi-Cevita covariant derivative has the following representation of the Christoffel symbols:
@@
The Levi-Cevita connection $\nabla$ is particularly suitable since we will see below that locally minimizing curves $\gamma$ of the energy (2.6) verify $\nabla_{\dot{\gamma}}\dot{\gamma}=0$ [112, Corrolary 6.7]. In other words, a local minimizing curve $\gamma$ has a vanishing second order derivative according to the geometry induced by the metric $\mathfrak{g}$. See Appendix A for an infinite dimensional extension of the covariant derivative which is relevant to analyze diffeomorphism groups or Wasserstein space geometry.
@@
With the covariant derivative, we can introduce the notion of parallel vector fields and the related concept of parallel transport [112, Chapter 4, p.60].
[…]
@@
Figure 4: Illustration of the parallel transport along the curve $\gamma_{x,v_{1}}$ (already represented in Figure 2).
The parallel transport of a vector along a curve is depicted in Figure 4. Finally, the concept of ‘straight lines’ in Euclidean spaces can be generalized to manifolds using geodesics. A geodesic $\gamma$ transports its own velocity vector in a parallel manner [112, Chapter 4, p.58].
[…]
@@
When using the affine invariant metric on $\mathcal{S}_{++}^{d}$, the geodesic $\gamma_{V,\Sigma}$ in direction $V\in\mathcal{S}^{d}$ with starting point $\Sigma\in\mathcal{S}_{++}^{d}$ is $t\mapsto\Sigma^{1/2}\exp(t\Sigma^{-1/2}V\Sigma^{-1/2})\Sigma^{1/2}$. In case of $\Sigma=\mathrm{Id}$, it reduces to the evaluation of the matrix exponential [144]. The parallel transport of a vector $W\in T_{\Sigma}\mathcal{S}^{d}$ at time $s=0$ to time $t$ along the geodesic $\gamma_{V,\Sigma}$ is $\mathfrak{T}_{\gamma_{V,\Sigma}(0)\rightarrow\gamma_{V,\Sigma}(t)}W=\exp(tV\Sigma^{-1}/2)W\exp(t\Sigma^{-1}V/2)$ [159, Lemma 3].
[…]
@@
[…]
Another way to characterize geodesics is through techniques from geometric mechanics. The second-order geodesic ODE (2.15) can be reformulated as a Hamiltonian system consisting of two first-order ODEs. To achieve this, we introduce the Hamiltonian for any $x,p\in\mathcal{M}\times T\mathcal{M}$,
@@
Based on any of these definitions, one can introduce the exponential map [112, Proposition 5.7], which “shoots” geodesics in prescribed directions as illustrated in Figure 2.
[…]
@@
This map is called the exponential map because of its link with the exponential on matrices when using the affine-invariant metric on $\mathcal{S}_{++}^{d}$ as detailed in Example 4.
In general, one can show that the exponential map is locally bijective, i.e. there exists an $\eta>0$, such that $\mathrm{Exp}_{p}\colon B_{\eta}(0)\to\mathrm{Exp}_{p}(B_{\eta}(0))$ is a bijection denoting by $B_{\eta}(0)$ the centered open ball of radius $\eta$. This follows from the local uniqueness of solutions of the geodesic equation. The image $U_{p}\coloneqq\mathrm{Exp}_{p}(B_{\eta}(0))$ is called a normal neighborhood [112, Lemma 5.10] of $p$ and, in this neighborhood, we can define the Riemannian logarithm as inverse of the exponential map:
[…]
@@
See Figure 5 for an illustration and Figure 6 for an application.
@@
Finally, to consider probability distributions and statistics on Riemannian manifold, we discuss how to integrate on them [151, Section II.5]. A metric $\mathfrak{g}$ introduces a natural volume form $\text{vol}_{\mathfrak{g}}$ on the manifold $\mathcal{M}$ through its determinant. That is, for a smooth function $f\colon\mathcal{M}\to\mathbb{R}$, we obtain the integral
@@
where the manifold is parametrized using $\psi\colon U\subset\mathbb{R}^{d}\to\mathcal{M}$. In case multiple functions are necessary to parametrize $\mathcal{M}$, one defines a partition of unity [151, Section I.2.1 and p.62] corresponding to the cover given by the images of the parametrizations and uses it to define the integral by combining the integrals defined for each image as above.
@@
The Euclidean structures are particularly easy to work with computationally, but they largely disregard the specific structure of SPD matrices, e.g. ¨it is hard to prevent nonpositive eigenvalues using it. In contrast, affine-invariant structures push symmetric matrices with nonpositive eigenvalues at an infinite distance, but incur high computational costs as it typically leads to an extensive use of matrix inverses, square roots, logarithms, and exponentials. Log-Euclidean metrics were introduced to achieve a compromise between these two properties. Through parametrizing $\mathcal{S}_{++}^{n}$ over the Euclidean space $\mathcal{S}^{n}$ of symmetric matrices computations become simply Euclidean in the logarithmic domain, limiting the number of necessary evaluations of the matrix logarithm and exponential yet the results are similar to the affine-invariant case [4]. Furthermore, the Log-Euclidean setting yields a Lie group structure in SPD matrices, i.e. turns it into a space that is both a manifold and a group where multiplication and inversion are smooth mappings. We will later discuss in Section 4.1 how to parametrize learnable metrics on SPD manifolds using these various structures.
@@
which is a Lie group, i.e. also equipped with a smooth manifold structure. It is one of the two connected components of the orthogonal group $O(n)\coloneqq\left\{Q\in\mathbb{R}^{n\times n}\mid Q^{T}Q=\mathrm{Id}_{n}\right\}$ which consists of all distance-preserving linear transformations of $\mathbb{R}^{n}$. The tangent space of $SO(n)$ (and also $O(n)$) at the identity is the Lie algebra of skew-symmetric matrices $T_{\mathrm{Id}}SO(n)=\mathfrak{so}(n)\coloneqq\{A\in\mathbb{R}^{n\times n}\mid A^{T}=-A\}$ and the tangent space at an arbitrary element $Q$ is given by the transformation of this space under $Q$, i.e. $T_{Q}SO(n)=Q\cdot T_{\mathrm{Id}}SO(n)\coloneqq\{QA\mid A\in\mathfrak{so}(n)\}$. Since these tangent spaces are also matrix spaces, to obtain a Riemannian metric one uses the Euclidean metric on matrices and obtains $\mathfrak{g}_{Q}(QA,QB)=\mathfrak{g}_{\mathrm{Id}}(A,B)=\frac{1}{2}\mathrm{tr}A^{T}B$ [^3] The factor $\frac{1}{2}$ is for convenience when defining metrics on the following matrix manifolds.. In dimensions two and three, $SO(n)$ consists of rotations around points or lines respectively and is therefore also called the rotation group. The (special) Euclidean group consists of its extension to all distance-preserving affine transformations, i.e. also considering translations: $SE(n)\coloneqq SO(n)\times\mathbb{R}^{n}$. Concerning applications, products of $SO(n)$ or $SE(n)$ are commonly used to model the rotation of joints, e.g. of skeleton models or rigs in computer animation [109] or of robotic arms [32, 16].
@@
That means it can be described as a further quotient manifold by $\mathrm{Gr}_{k}(\mathbb{R}^{n})\cong\mathrm{St}_{k}(\mathbb{R}^{n})/O(k)$, where the action of $O(k)$ is by change of the orthonormal basis of the subspace. We focus here on another description, namely via projection matrices. For this one identifies a linear subspace with the unique orthogonal projection onto it and obtains

### v2

# A Review on Riemannian Metric Learning: Closer to You than You Imagine

@@ Abstract
To recount the story from the beginning, let us begin with the principles of metric learning. Metric learning [22] is an integral, though often implicit, aspect of machine learning. Whether consciously or not, most practitioners in the field are already engaging with it—or will eventually. The concept has a long history, dating back to early methods like the k-nearest neighbors (k-NN) classifier [45, 169, 67]. As the term ‘nearest’ suggests, a measure of distance is at the heart of k-NN algorithms as depicted in Figure˜1. Without prior knowledge, the distance is often selected as the Euclidean distance. However, this choice is uninformative of the input and output distribution. That is why Atkeson et al. [11] suggested to learn the distance by learning a symmetric positive definite matrix $G\in\mathbb{R}^{d\times d}$ such that $\mathbf{d}_{G}(x,y)=\sqrt{(y-x)^{\top}G(y-x)}$. It turns out that learning $G=L^{\top}L$ is equivalent to learning a linear transformation $x\mapsto Lx$ of the input space as $\mathbf{d}_{G}=\mathbf{d}(L\cdot,L\cdot)$. The matrix $G$ can be learned, for example, in an unsupervised manner using Principal Component Analysis (PCA) or in a supervised manner using Linear Discriminant Analysis (LDA). Learning $G$ is valuable for two key reasons: First, it assigns global relative weights to the different coordinates when computing distances. Second, it allows us to account for global correlations in the dataset. However, $G$ needs to vary to account for heterogeneous local geometries within the dataset as shown in Figure˜1. This linear framework was then extended with non-linear transformations $x\mapsto\phi(x)$ by using kernels or neural networks [205, 55, 69] such that the distance is $\mathbf{d}(\phi(\cdot),\phi(\cdot))$, which reduces to learning a feature function $\phi$ as already pointed out.
@@
Going one step further, metric learning can be embedded in a more general framework called Riemannian metric learning [147], in the same way that Galilean relativity is embedded by General relativity [60]. In Riemannian metric learning, the input space $(\mathcal{X},{\mathfrak{g}})$ is a Riemannian manifold, i.e. a kind of smooth geometric shape (see Figure˜2) with a Riemannian metric ${\mathfrak{g}}\colon\mathcal{X}\to\mathbb{R}^{d\times d}$ that allows to locally measure lengths and angles via a matrix field. From this local geometric structure, one can define a global distance between any $x,y\in\mathcal{X}$ as
@@
Minimizing curves of (1.1) are called metric geodesics and can be understood as generalizations of straight lines to manifolds as illustrated in Figure˜1. If the Riemannian metric is constant and the manifold is the Euclidean space, then geodesics are straight lines and we recover the distance metric learning framework from above. However, Riemannian metric learning is not just a generalization of distance metric learning with non-linear transformations; it also gives new mathematical objects useful for modeling. For instance, geodesics can express trajectories for longitudinal modeling and interpolation [126, 112, 103], parallel transport (defined later) can be used in transfer learning [179, 76], and the metric volume can be used for modeling probability distribution and sampling [47, 50]. All these mathematical objects are related to each other and have been extensively studied in Riemannian Geometry [53, 171]. Riemannian metric learning leverages the wealth of this literature to offer both flexible and interpretable modeling tools.
@@
First, we recall preliminary concepts on the essential tools and properties of Riemannian geometry in Section˜2; readers familiar with these notions may safely skip this section. At the same time, we introduce the main manifolds relevant to the data used in Riemannian metric learning in Section˜2.1, along with the manifold of Riemannian metrics, to provide a broader perspective on the subject. Second, the formal optimization problem associated with Riemannian metric learning is presented in Section˜3.1 to define the topic rigorously. This is followed by an overview of its numerous applications in Section˜3.2, highlighting the practical benefits and key techniques of Riemannian formalism. Third, we discuss how to perform Riemannian metric learning in Section˜4, classifying the methods based on their metric parameterization (Section˜4.1) and associated computational approaches (Section˜4.1.3), their objectives (Section˜4.2), and their optimization strategies (Section˜4.3). This classification helps clarify the various technical challenges that arise when designing Riemannian metric learning methods. Finally, before concluding in Section˜6, we outline key research directions in Riemannian metric learning in Section˜5, which may inspire both theorists and practitioners in the applied mathematics community.
@@ Acknowledgments
We denote by $\mathcal{P}(\mathsf{X})$ the power set of a set $\mathsf{X}$, integer ranges by $k:l]=\{k,\ldots,l\}\subset\mathcal{P}(\mathbb{Z})$ and $[l]=[1:l]$ with $k,l\in\mathbb{N}$. The canonical scalar product of two vectors $x=(x_{i})_{i\in[d]},y=(y_{i})_{i\in[d]}\in\mathbb{R}^{d}$ is denoted by $x^{\top}y=\sum_{i=1}^{d}x_{i}y_{i}$. The Euclidean norm is denoted by $||\cdot||=\sqrt{\cdot^{\top}\cdot}$. We denote by $\mathrm{C}^{k}(\mathsf{U},\mathbb{R}^{d})$ the set of all $k$ times continuously differentiable functions from $\mathsf{U}$ to $\mathbb{R}^{d}$. When $\mathsf{U}\subset\mathbb{R}$, we denote by $\dot{\gamma}$ the time derivative of $\gamma\in\mathrm{C}^{1}(\mathsf{U},\mathbb{R}^{d})$. The spaces of square symmetric and symmetric positive definite matrices are denoted by $\mathcal{S}^{d}=\{A\in\mathbb{R}^{d\times d}\,:\,A^{\top}=A\}$ and $\mathcal{S}^{d}_{++}=\{A\in\mathcal{S}^{d}\,:x^{\top}Ax>0\,\text{for any}\,x\in\mathbb{R}^{d}\setminus\{0\}\}$, respectively. The square identity matrix in dimension $d\times d$ is denoted by $\mathrm{I}_{d}$. The determinant and the trace of a matrix $A\in\mathbb{R}^{d\times d}$ are denoted by $\mathrm{tr}(A)$ and $\det(A)$ respectively. For any $(x_{i})_{i\in[m]}\in\mathbb{R}^{m}$, $\operatorname{diag}((x_{i})_{i\in[m]})$ denotes the diagonal matrix whose diagonal elements are $(x_{i})_{i\in[m]}$, arranged from the top-left. For any $f\in\mathrm{C}^{1}(\mathsf{U},\mathbb{R}^{m})$ with $\mathsf{U}\subset\mathbb{R}^{d}$, $\mathop{}\!\mathrm{d}f$ is the differential of $f$. For any $f\in\mathrm{C}^{1}(\mathsf{U},\mathbb{R})$ with $\mathsf{U}\subset\mathbb{R}^{d}$, $\nabla f$ is the gradient of $f$, but can also denotes the covariate derivative depending on the context. For any $j\in[d]$ and any $\psi\in\mathrm{C}^{1}(\mathsf{U},\mathbb{R}^{m})$ with $\mathsf{U}\subset\mathbb{R}^{d}$, $\partial_{j}\psi\in\mathbb{R}^{m}$ denotes the directional directive of $\psi$ in direction $e_{j}$, where $e_{j}$ is the $j$-th element of the canonical basis. When a random variable $X\in E$ follows a probability distribution $P$ on a measurable space $(E,\mathcal{B})$, we write $X\sim P$. $\mathcal{U}([0,1])$ denotes the uniform probability distribution on $[0,1]$. If $P_{1}$ and $P_{2}$ are two probability distribution on $(E,\mathcal{B}_{1})$ and $(F,\mathcal{B}_{2})$ respectively, then $P_{1}\otimes P_{2}$ denotes their product measure.
@@
A differentiable manifold $\mathcal{M}\subset\mathbb{R}^{d_{a}}$ is formally defined in Appendix˜A or in [115, Chapter 2], where $d_{a}$ denotes the dimension of the ambient space. For a first reading, and to simplify notation [^1] All concepts introduced below are compatible with changes of parametrization in overlapping regions and can therefore be generalized to manifolds requiring multiple parametrizations., a differentiable manifold of dimension $d<\infty$ can be understood as the image of a global, surjective, and smooth parametrization $\psi\colon U\subset\mathbb{R}^{d}\to\mathcal{M}$, where $U$ is an open subset of $\mathbb{R}^{d}$. In general, we would need more than one parametrization and all concepts presented hereafter generalize to this scenario. Readers unfamiliar with this concept may safely think of a differentiable manifold as a smoothly curved, nonlinear object, such as a curved surface in three-dimensional space as depicted in Figure˜2. Note that in all the following a function $f\colon\mathcal{M}\to\mathcal{Y}\subset\mathbb{R}^{m}$ is called smooth or infinitely differentiable if $f\circ\psi\colon U\to\mathcal{Y}$ is so, which offers a valid definition since $U$ is an open set of $\mathbb{R}^{d}$.
@@
The dimension of the manifold is thus given by the dimension of possible local displacements in each point. In the case of two-dimensional surfaces, the tangent space is simply the plane that is tangent to the surface at a given point as illustrated in Figure˜2.
@@
More manifold examples and insights related to computations and modeling will be given in Section˜2.1.
@@
This path length is invariant to reparametrization of the curve [115, Lemma 6.1]. Furthermore, the path energy is defined as
@@
and equality holds if and only if ${\mathfrak{g}}_{\gamma(t)}(\dot{\gamma}(t),\dot{\gamma}(t))$ is constant along the path, which is the case for geodesics as we will see later. Given the length of paths, we can define the distance between two arbitrary points as the shortest possible length of a path connecting them, which leads to the definition of a Riemannian distance [115, Lemma 6.2]:
Definition 6 (Riemannian distance).
@@
Shortest paths can have different shapes depending on the metric as illustrated in Figure˜3.
@@
In order to find optimality constraints related to (2.6), we need to compute variations of $\mathcal{L}$ with respect to $\gamma$ which thus implies second order derivatives $\ddot{\gamma}$. However, defining second order derivatives via a difference quotient $\lim_{dt\to 0}(\dot{\gamma}(t+dt)-\dot{\gamma}(t))/dt$ is not straightforward on manifolds. The vectors $\dot{\gamma}(t+dt)\in T_{\gamma(t+dt)}\mathcal{M}$ and $\dot{\gamma}(t)\in T_{\gamma(t)}\mathcal{M}$ belong to two different spaces as depicted in Figure˜2 and thus cannot we cannot subtract them a priori. Thus, we introduce a way to do so next.
@@
We define a covariant derivative $\nabla$ as a map $C^{\infty}(\mathcal{M},T\mathcal{M})\times C^{\infty}(\mathcal{M},T\mathcal{M})\to C^{\infty}(\mathcal{M},T\mathcal{M})$ which satisfies intuitive derivation rules as exposed in the following example or in Appendix˜A given in Appendix˜A or in [115, Chapter 4, p.50]. Intuitively, $\nabla_{W}V$ describes how the vector field $V$ changes when moving in direction $W$. In this way, a covariant derivative $\nabla$ of a vector field $V=\sum_{i=1}^{d}V_{i}\partial_{i}\psi$ in direction of $W=\sum_{i=1}^{d}W_{i}\partial_{i}\psi$ denoted by $\nabla_{W}V$ can be decomposed as
@@
for any $x\in\mathcal{M}$ and $Z,V,W\in C^{\infty}(\mathcal{M})$. This defines the Levi-Cevita covariant derivative $\nabla$ also called the Levi-Cevita connection [115, Theorem 5.4]. The following property gives a concrete expression.
Proposition 9.
Let ${\mathfrak{g}}^{-1}=({\mathfrak{g}}^{ij})_{(i,j)\in [d]^{2}}$ be the inverse of $\mathfrak{g}$} as before. Then, the Levi-Cevita covariant derivative has the following representation of the Christoffel symbols:
@@
The Levi-Cevita connection $\nabla$ is particularly suitable since we will see below that locally minimizing curves $\gamma$ of the energy (2.6) verify $\nabla_{\dot{\gamma}}\dot{\gamma}=0$ [115, Corrolary 6.7]. In other words, a local minimizing curve $\gamma$ has a vanishing second order derivative according to the geometry induced by the metric $\mathfrak{g}$}. See Appendix˜A for an infinite dimensional extension of the covariant derivative which is relevant to analyze diffeomorphism groups or Wasserstein space geometry.
@@
With the covariant derivative, we can introduce the notion of parallel vector fields and the related concept of parallel transport [115, Chapter 4, p.60].
Proposition 10 (Parallel transport).
@@
Figure 4: Illustration of the parallel transport along the curve $\gamma_{x,v_{1}}$ (already represented in Figure˜2).
The parallel transport of a vector along a curve is depicted in Figure˜4. Finally, the concept of ‘straight lines’ in Euclidean spaces can be generalized to manifolds using geodesics. A geodesic $\gamma$ transports its own velocity vector in a parallel manner [115, Chapter 4, p.58].
Definition 11 (Geodesic).
@@
When using the affine invariant metric on $\mathcal{S}_{++}^{d}$, the geodesic $\gamma_{V,\Sigma}$ in direction $V\in\mathcal{S}^{d}$ with starting point $\Sigma\in\mathcal{S}_{++}^{d}$ is $t\mapsto\Sigma^{1/2}\exp(t\Sigma^{-1/2}V\Sigma^{-1/2})\Sigma^{1/2}$. In case of $\Sigma=\mathrm{Id}$, it reduces to the evaluation of the matrix exponential [148]. The parallel transport of a vector $W\in T_{\Sigma}\mathcal{S}^{d}$ at time $s=0$ to time $t$ along the geodesic $\gamma_{V,\Sigma}$ is $\mathfrak{T}_{\gamma_{V,\Sigma}(0)\rightarrow\gamma_{V,\Sigma}(t)}W=\exp(tV\Sigma^{-1}/2)W\exp(t\Sigma^{-1}V/2)$ [164, Lemma 3].
Remark 13 (Variational definition).
@@
Remark 14 (Hamiltonian flow).
Another way to characterize geodesics is via techniques from geometric mechanics. The second-order geodesic ODE (2.15) can be reformulated as a Hamiltonian system consisting of two first-order ODEs. To achieve this, we introduce the Hamiltonian for any $x,p\in\mathcal{M}\times T\mathcal{M}$,
@@
Based on any of these definitions, one can introduce the exponential map [115, Proposition 5.7], which “shoots” geodesics in prescribed directions as illustrated in Figure˜2.
Definition 15 (Exponential map).
@@
This map is called the exponential map because of its link with the exponential on matrices when using the affine-invariant metric on $\mathcal{S}_{++}^{d}$ as detailed in Example˜12.
In general, one can show that the exponential map is locally bijective, i.e. there exists an $\eta>0$, such that $\mathrm{Exp}_{p}\colon B_{\eta}(0)\to\mathrm{Exp}_{p}(B_{\eta}(0))$ is a bijection denoting by $B_{\eta}(0)$ the centered open ball of radius $\eta$. This follows from the local uniqueness of solutions of the geodesic equation. The image $U_{p}\coloneqq\mathrm{Exp}_{p}(B_{\eta}(0))$ is called a normal neighborhood [115, Lemma 5.10] of $p$ and, in this neighborhood, we can define the Riemannian logarithm as inverse of the exponential map:
Definition 16 (Logarithm).
@@
See Figure˜5 for an illustration and Figure˜6 for an application.
@@
Finally, to consider probability distributions and statistics on Riemannian manifold, we discuss how to integrate on them [156, Section II.5]. A metric $\mathfrak{g}$} introduces a natural volume form $\text{vol}_{\mathfrak{g}}$ on the manifold $\mathcal{M}$ through its determinant. That is, for a smooth function $f\colon\mathcal{M}\to\mathbb{R}$, we obtain the integral
@@
where the manifold is parametrized using $\psi\colon U\subset\mathbb{R}^{d}\to\mathcal{M}$. In case multiple functions are necessary to parametrize $\mathcal{M}$, one defines a partition of unity [156, Section I.2.1 and p.62] corresponding to the cover given by the images of the parametrizations and uses it to define the integral by combining the integrals defined for each image as above.
@@
The Euclidean structures are particularly easy to work with computationally, but they largely disregard the specific structure of SPD matrices, e.g. it is hard to prevent nonpositive eigenvalues using it. In contrast, affine-invariant structures push symmetric matrices with nonpositive eigenvalues at an infinite distance, but incur high computational costs as it typically leads to an extensive use of matrix inverses, square roots, logarithms, and exponentials. Log-Euclidean metrics were introduced to achieve a compromise between these two properties. Through parametrizing $\mathcal{S}_{++}^{n}$ over the Euclidean space $\mathcal{S}^{n}$ of symmetric matrices computations become simply Euclidean in the logarithmic domain, limiting the number of necessary evaluations of the matrix logarithm and exponential yet the results are similar to the affine-invariant case [5]. Furthermore, the Log-Euclidean setting yields a Lie group structure in SPD matrices, i.e. turns it into a space that is both a manifold and a group where multiplication and inversion are smooth mappings. We will later discuss in Section˜4.1 how to parametrize learnable metrics on SPD manifolds using these various structures.
@@
which is a Lie group, i.e. also equipped with a smooth manifold structure. It is one of the two connected components of the more general orthogonal group $O(n)\coloneqq\left\{Q\in\mathbb{R}^{n\times n}\mid Q^{T}Q=\mathrm{Id}_{n}\right\}$ which consists of all distance-preserving linear transformations of $\mathbb{R}^{n}$. The tangent space of $SO(n)$ (and also $O(n)$) at the identity is the Lie algebra of skew-symmetric matrices $T_{\mathrm{Id}}SO(n)=\mathfrak{so}(n)\coloneqq\{A\in\mathbb{R}^{n\times n}\mid A^{T}=-A\}$ and the tangent space at an arbitrary element $Q$ is given by the transformation of this space under $Q$, i.e. $T_{Q}SO(n)=Q\cdot T_{\mathrm{Id}}SO(n)\coloneqq\{QA\mid A\in\mathfrak{so}(n)\}$. Since these tangent spaces are also matrix spaces, to obtain a Riemannian metric one uses the Euclidean metric on matrices and obtains ${\mathfrak{g}}_{Q}(QA,QB)={\mathfrak{g}}_{\mathrm{Id}}(A,B)=\frac{1}{2}\mathrm{tr}A^{T}B$ [^3] The factor $\frac{1}{2}$ is for convenience when defining metrics on the following matrix manifolds.. In dimensions two and three, $SO(n)$ consists of rotations around points or lines respectively and is therefore also called the rotation group. The (special) Euclidean group consists of its extension to all distance-preserving affine transformations, i.e. also considering translations: $SE(n)\coloneqq SO(n)\times\mathbb{R}^{n}$. Concerning applications, products of $SO(n)$ or $SE(n)$ are commonly used to model the rotation of joints, e.g. of skeleton models or rigs in computer animation [112] or of robotic arms [35, 18].
@@
That means it can be described as a further quotient manifold by observing $\mathrm{Gr}_{k}(\mathbb{R}^{n})\cong\mathrm{St}_{k}(\mathbb{R}^{n})/O(k)$, where the action of $O(k)$ is by change of the orthonormal basis of the subspace. We focus here on another description, namely via projection matrices. For this one identifies a linear subspace with the unique orthogonal projection onto it and obtains